OpenStudy (anonymous):

Anton joined a golf club two years ago. he pays annual fee of $895 and a fee of $30 each time he plays a game of golf. The function below can be used to calculate the totql yearly golfing fee, f(g)=895+30g where g represents the number of times he played golf during the year. Last year he paid $2,065 as a total golfing fee. for how many games did he pay a greens fee? please explain.

OpenStudy (anonymous):

required 2065 = 895 +30g g= (2065-895)/30

OpenStudy (anonymous):

39